OpenStudy (anonymous):

A circle whose center is at (1, -3) passes through (7, 5). What is the length of the radius of the circle?

OpenStudy (anonymous):

just find the distance between the points (1, -3) and (7, 5).

OpenStudy (anonymous):

reqd distance is sqrt((7-1)^2 + (5-(-3))^2)=sqrt(36+64)=sqrt(100)=10